OpenStudy (anonymous):

secxcsc^2x=2secx how to solve equation in the given interval

OpenStudy (anonymous):

interval is a set of real numbers with the property that any number that lies between two numbers in the set is also included in the set. For example, the set of all numbers x satisfying 0 ≤ x ≤ 1 is an interval which contains 0 and 1, as well as all numbers between them.

OpenStudy (anonymous):

\[\sec x \csc ^2x-2\sec x=0,\sec x \left( \csc ^2x-2 \right)=0\] \[Either ~\sec x=0,\cos x=\infty ,rejected~as~\left| \cos x \right|\le 1\]

OpenStudy (anonymous):

what's the given interval?

OpenStudy (anonymous):

oh sorry! 0 less than or equal to x les than 2pi

OpenStudy (anonymous):

0<=x<2pi

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[or \csc ^2x=2,\sin ^2x=\frac{ 1 }{2 },2\sin ^2x=1,1-2\sin ^2x=0,\cos 2x=0\] find x

OpenStudy (anonymous):

can you solve?

OpenStudy (anonymous):

nope:/

OpenStudy (anonymous):

\[\cos 2x=0=\cos \left( 2n+1 \right)\frac{ \pi }{2 },n=0,1,2,3\] \[2x=\left( 2n+1 \right)\frac{ \pi }{2 },x=?\]

OpenStudy (cwrw238):

or you can convert each trig ratio to cos and sin and get and equation in cos x

OpenStudy (cwrw238):